Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

10 light bulbs

Expert replies
by massi2884 » Mon Apr 23, 2012 11:18 am
I know it's not among the answer choices, but could you please tell me what's wrong with thinking that the result should be
0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 = 10*0.06 = 0.6 ?

A string of 10 light bulbs is wired in such a way that if any individual light bulb fails, the entire string fails. If for each individual light bulb the probability of failing during time period T is 0.06, what is the probability that the string of light bulbs will fail during the time period T?

(A) 0.06
(B) (0.06)^10
(C) 1 - (0.06)^10
(D) (0.94)^10
(E) 1 - (0.94)^10

OA E
Join the discussion
Source: — Problem Solving |

by Shalabh's Quants » Mon Apr 23, 2012 12:08 pm
massi2884 wrote:I know it's not among the answer choices, but could you please tell me what's wrong with thinking that the result should be
0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 = 10*0.06 = 0.6 ?

A string of 10 light bulbs is wired in such a way that if any individual light bulb fails, the entire string fails. If for each individual light bulb the probability of failing during time period T is 0.06, what is the probability that the string of light bulbs will fail during the time period T?

(A) 0.06
(B) (0.06)^10
(C) 1 - (0.06)^10
(D) (0.94)^10
(E) 1 - (0.94)^10

OA E
Probability of String does not fail = (1-0.06)^10; All 10 bulbs are glowing.

Hence, to fail; prob. = 1- prob. of success = 1 - (1-0.06)^10 = 1-0.94^10.
Shalabh Jain,
e-GMAT Instructor
Join the discussion

by Anurag@Gurome » Mon Apr 23, 2012 7:48 pm
massi2884 wrote:I know it's not among the answer choices, but could you please tell me what's wrong with thinking that the result should be
0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 = 10*0.06 = 0.6 ?

A string of 10 light bulbs is wired in such a way that if any individual light bulb fails, the entire string fails. If for each individual light bulb the probability of failing during time period T is 0.06, what is the probability that the string of light bulbs will fail during the time period T?

(A) 0.06
(B) (0.06)^10
(C) 1 - (0.06)^10
(D) (0.94)^10
(E) 1 - (0.94)^10

OA E
Probability of 1 light bulb failing during time period T = 0.06
So, probability of 1 light bulb NOT failing during time period T = 1 - 0.06 = 0.94

For the string to be successful, all the light bulbs should pass because if any individual light bulb fails, the entire string fails.
Probability of 10 light bulbs NOT failing = (0.94)^10
Probability of 10 light bulbs failing = 1 - Probability of 10 light bulbs NOT failing = 1 - (0.94)^10

The correct answer is E.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by Mr Smith » Tue Apr 24, 2012 3:27 am
Yes you should always ask such question to gain a better understanding of all concepts.

Let the event of the first light bulb failing be E1, for the 2nd be E2 and so on.
OK what we need is the Union of the events E1, E2... E10 i.e., P(E1 OR E2 OR E3 OR .... OR E10)
From sets we know that n(A union B) = n(A) + n(B) - n(A intersection B)..
Extending to probability P(E1 OR E2) = P(E1) + P(E2) - P(E1 AND E2)
Such a formula can also be derived for 10 events. But the Key here is that its not simply equal to the sum of n(A) and n(B). When you simply add the probability you are also adding the occurance of events like {E1 AND E2} etc. twice.
Now if you want to find this union, one way is to derive such a formula for 10 events and then calculate the probability using it which is very tedious.

So a better way is to find the probability of the P(NOT E1 AND NOT E2...... AND NOT E10) and subtract it from 1 as aptly demonstrated by others.
P(E1 OR E2 OR E3 OR .... OR E10) = 1-P(NOT E1 AND NOT E2...... AND NOT E10)
= 1-P(NOT E1)*P(NOT E2)*....*P(NOT E10) ---- since independent events
= 1-[(1-P(E1))* (1-P(E2)) *(1-P(E3))* ..... *(1-P(E10))]
Kindly Use Thanks Button as liberally as you receive replies. :)

Mr. Smith
Join the discussion